OpenStudy (anonymous):

f(X) =2x^2-3x+1 find f(-1) F(1/4) f(0.3) given the function F(x) find the following; =2x^2-3x+1 find f(-1) F(1/4) f(0.3)

OpenStudy (anonymous):

f(-1) = 2(-1)^2 - 3(-1) + 1 = 2 + 3 + 1 = 6 f(.3) = 2(.3)^2 - 3(.3) + 1 = .18 - .9 + 1 = 1.72 F(.25) = f'(.25) = 4(.25) - 3 = -2

OpenStudy (anonymous):

in the first one isn't 2(-1)^2= -4?